# Finding the Fairest Voting System Using Likelihood Analysis

### v2

@@ 6.2 Pareto Unanimity (U)
In the following, a Pareto pair is defined to be a pair of alternatives where all voters prefer one alternative in the pair over the other.
Only Group 1 systems can theoretically violate U. Even for these systems, U is nearly always satisfied. In elections with at least 30 voters, all systems have a 0% probability of violating U. U is almost always satisfied because encountering a Pareto pair is very rare, especially as the number of voters increases. Table 4 shows the frequency with which Instant Runoff violates U.
Table 4: Percentage Unanimity Violations Instant Runoff
(a) IC Data
Number of voters Number of alternatives
@@
1 100.00 100.00 100.00 100.00
2 16.23 58.00 86.12 96.61
3 2.56 17.38 41.34 62.30
4 0.43 4.59 13.93 28.63
5 0.05 1.14 4.58 10.87
10 0.00 0.00 0.00 0.07
30 0.00 0.00 0.00 0.00
50 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
1 100.00 100.00 100.00 100.00
2 28.05 59.94 86.13 96.56
3 11.02 21.64 41.69 63.20
4 4.42 7.96 15.58 28.35
5 2.32 2.83 5.00 10.14
10 0.19 0.06 0.04 0.05
30 0.01 0.00 0.00 0.00
50 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
[…]
Table 5 shows U violations for elections with four alternatives using IAC for Group 1 systems. Figure 3 is a graph of the data.
Table 5: Percentage U Violations 4 alternatives IAC
[…]
1 2 3 4 5 10 30 50 100
Plurality/InstantRunoff/TopTwo 100.00 59.94 21.64 7.96 2.83 0.06 0 0 0
Minimax 100.00 44.53 12.93 4.04 0.98 0.01 0 0 0
Figure 3: U Violations in elections with four alternatives (IAC)
While unanimity does not pose issues in large elections, for election with few voters and many candidates, this criterion is violated often.
@@ 6.3 Transitivity (TR)
Pairwise Majority is the only system studied that violates transitivity. Table 6 shows the percentage TR violations for Pairwise Majority. Figure 4 is a 3D plot of the IC data.
@@
    As the number of alternatives increases, the transitivity violations increase.
@@
[…]
@@
30000 6.45 24.06 46.92 66.98
50000 6.34 23.90 46.49 67.21
Figure 4: TR Violations in Pairwise Majority elections (IC)
[…]
6.4 Joint Satisfaction of the Conditions
Simulations were run to find the percentage of times a voting system violates at least one criteria. The order in which criteria were tested was U, IIA, and TR. The other criterion (ND, UD) were not simulated as no system violates ND or UD. Joint-Satisfaction (JS) Frequencies, defined as the percentage of times systems satisfied all criteria simultaneously, serve as an indication of fairness of a voting system.
For every system that violates IIA, the JS Frequencies almost exactly align with the likelihood of satisfying IIA. In elections with more than three alternatives and two voters, no system had a JS Frequency of 100%, which is consistent with Arrow’s Impossibility Theorem. Table 7 shows the results for Baldwin. All other data tables can be found in Appendix A.
Table 7: JS Frequencies (%) for Baldwin
@@ 6.4 Joint Satisfaction of the Conditions
10 5.16 0.01 0.00 0.00
30 0.10 0.00 0.00 0.00
50 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
5000 0.00 0.00 0.00 0.00
10000 0.00 0.00 0.00 0.00
30000 0.00 0.00 0.00 0.00
50000 0.00 0.00 0.00 0.00
(b) IAC Data
Number of voters Number of alternatives
3       4       5 6
10 25.16 0.25 0.01 0.00
30 16.87 0.27 0.00 0.00
50 16.79 0.28 0.00 0.00
100 17.14 0.16 0.00 0.00
1000 17.27 0.22 0.00 0.00
5000 16.63 0.27 0.00 0.00
10000 16.68 0.29 0.00 0.00
30000 17.07 0.22 0.00 0.00
50000 16.88 0.33 0.00 0.00
Table 8 shows a comparison of all the voting systems for 3 alternative elections under IAC and Figure 5 is a graph of the data.
Table 8: Joint-Satisfaction Frequencies for 3 alternatives (IAC)
Voting System Number of Voters
10 100 1000 5000 10000 50000
Group 1 0.00 0.00 0.00 0.00 0.00 0.00
Copeland 1.90 0.02 0.00 0.02 0.03 0.02
Dowdall 7.41 3.16 3.13 3.11 2.97 3.10
Borda Count 25.16 16.15 16.00 15.48 15.37 15.75
Coombs 25.16 14.50 14.20 13.90 13.40 14.02
Baldwin 25.16 17.14 17.27 16.63 16.68 16.88
Pairwise Majority 84.50 92.77 93.54 93.76 93.67 93.66
Figure 5: JS Frequencies in elections with three alternatives (IAC)
Figure 6: JS Frequencies for Black, Borda, Baldwin, Coombs
The data shows that Pairwise Majority is the most likely to satisfy all the criteria jointly for elections with three alternatives. As the number of voters increase in Pairwise Majority, the JS Frequencies increase, while for other systems the JS Frequencies either decrease or remain approximately constant. This is due to the fact that as the number of voters increase, transitivity violations become more rare, while, in general, IIA violations become more common.
There are also clear differences in performance in the other systems in three-alternative elections. Of the systems that satisfy transitivity, Borda Count, Baldwin, and Coombs have the highest JS Frequencies. Figure 6 shows a zoomed in version of Figure 5 for these three systems. For Borda, Baldwin, and Coombs the JS Frequencies also slightly increase as the number of voters becomes very large. This is because the IIA violations in these systems tend to marginally decrease as the number of voters increases from 30 to 50,000.
However, as the number of alternatives increases, the JS Frequencies decrease for every system. Figure 7 shows this trend clearly. In fact, for all systems excluding Pairwise Majority, the JS Frequencies for elections with at least 10 voters and 4 alternatives are less than 0.5%. This is because IIA is almost always violated in elections with at least four alternatives. Even for Pairwise Majority, the JS Frequencies sharply decrease as the number of alternatives increase, falling from 93.66% for three alternatives elections to 32.80% for six alternatives elections.
Figure 7: JS Frequencies in elections with 50,000 voters (IAC)
The conclusion for three-alternative elections is consistent with literature. In Dougherty and Heckelman (2020), the likelihood of violating at least one of Arrow’s conditions is computed. Their study agrees that Pairwise Majority has the greatest JS Frequencies. However, due to differences in the IIA data as discussed in Section 7, the data for Copeland, Coombs, Dowdall, Borda, and Baldwin does not match exactly. Most systems have lower JS Frequencies than in the 2020 study.
7 Remarks about Differences in IIA Results
The IIA algorithm used in this paper is different than that of Dougherty and Heckelman (2020) and we describe those differences here. We start by describing our algorithm.
As IIA is an inter-profile condition, to check if a voting system violates IIA, multiple preference schedules have to be compared. Our IIA algorithm loops through all possible pairs of alternative. For each alternative pair $A$ and $B$, the following test is done: First, find the societal preference between $A$ and $B$. Then, let $j$ equal the number of voters who prefer $A$ over $B$ and $k$ the number of voters who prefer $B$ over $A$. Then randomly generate 500 additional preference schedules in which $j$ voters still prefer $A$ over $B$ and $k$ voters prefer $B$ over $A$. The number 500 was chosen taking various factors into consideration including computational resources and program run time. If any of these new preference schedules produce a different societal preference between $A$ and $B$ than the original, an IIA violation occurs. This is repeated for every possible pair of alternatives.
This algorithm can miss violations, which is more likely when the number of voters is high. As the number of voters increases, there are more possible preference schedules where $j$ voters prefer $A$ over $B$ and $k$ voters vote $B$ over $A$. Thus, in the 500 new preference schedules generated, there may be no preference schedule that has a different societal order, even if one exists. Checking every single preference is impractical since there are $(\frac{c!}{2})^{v}$ possible schedules such that each voter maintains their relative preference between a pair, where $c$ is the number of alternatives and $v$ is the number of voters.
In Dougherty and Heckelman (2020), a different IIA algorithm was used. Our algorithm finds substantially more IIA violations that the algorithm in Dougherty and Heckelman (2020), in particular with the IAC data. For Dowdall, Coombs, Copeland, and Baldwin our algorithm reports a greater percentage IIA violations. For Coombs and Copeland, the difference in the IAC data is around 20 percentage points.
7.0.1 Disagreement in IIA Conclusions under Copeland
2 0 0 1 0 0
1st A A B B C C
2nd B C A C A B
3rd C B C A B A
Under Copeland, the societal preference order is $A\succ B\succ C$. Now, consider the pair of alternatives A and B. Suppose that one voter who has preference $A\succ_{i}B\succ_{i}C$ changes their preference to $C\succ_{i}A\succ_{i}B$. Now the preference schedule is:
1 0 0 1 1 0
1st A A B B C C
2nd B C A C A B
3rd C B C A B A
Dougherty and Heckelman (2020) also state that Copeland violates IIA at a greater frequency for even number of voters $N$ than for the adjacent odd values $N-1$ and $N+1$. They claim that this is the case as there are more ties for elections with an even number of voters. While ties do make it easier to violate IIA, there are many cases for four voter elections where there are no ties and IIA is satisfied. Our analysis shows that for even number of voters, IIA violations in Copeland actually occur less frequently. For instance in three-alternative IC elections, Copeland violates IIA 97.35% of the time for three voters and only 73.63% of the time for four voters under IC.
Another, possibly related, disagreement was found while using their algorithm. For the purpose of comparing algorithms, we coded their IIA algorithm in our Python model for testing. Using their algorithm, we obtained the exact same data table that they did for both IC and IAC, with the exception of Copeland. For Copeland elections with three alternatives and at least four voters (in elections with three voters, the data is in agreement), their reported values are higher than what their algorithm produces. For instance, in elections with four voters and three alternatives under IC, simulations using their algorithm show a 73.7% chance of violating IIA, while the value reported in their paper is 88.9%. Notably, our IIA tables show 73.6% violations for identical elections.
This over-counting is also seen for IAC elections with more than four voters. For elections with 30 voters and three alternatives under IAC, our implementation of their algorithm yields 72% IIA violations, but the reported value is 78.5%. Note that as the number of voters increases, the extent of over-counting decreases. For IC, their data reaches 100% violations, and for IAC the difference between their values and our implementation of their algorithm becomes almost 0. The authors suspect an error in their Copeland method for greater than four voters.
8 Summary of Results
After comparing all systems, Pairwise Majority is the fairest system as it has the greatest likelihood of jointly satisfying all social choice criterion in elections with at least three alternatives and two voters. The other systems perform very poorly in comparison, primarily due to the extremely high frequency of IIA violations. Of the systems that satisfy transitivity, Baldwin, Borda, and Coombs perform the best, but have JS Frequencies of less than 17%.
Of all systems considered, Group 1 systems perform the worst. They violate IIA 100% of the time and as such, always have a 0% probability of jointly satisfying all criterion in elections with at least three alternatives.
Another key observation is that as the number of alternatives increases, the frequency with which voting systems violate criteria generally increases. For IIA, TR, and U, most systems increase in percentage violations as the number of alternatives increase. Thus, in elections with more alternatives, the JS Frequencies are significantly lower.
In the end, in alignment with Arrow’s Theorem, no voting system is perfect. Each method gives rise to results which are arguably undesirable from the point of view of a society. However, not all voting systems are equally flawed. While some systems always violate social choice criteria, there are systems with a high likelihood of satisfying all criteria. Finding and implementing such systems can help achieve fairer and more representative electoral outcomes.
9 Declarations
Data Availability Statement The raw data generated during this study is available upon request. All relevant processed data is included within this article and the appendix.
[…]
Appendix A Data Tables
A.1 IIA
Table 9: Percentage IIA Violations Plurality and Instant Runoff
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
3 100.00 100.00 100.00 100.00
4 100.00 100.00 100.00 100.00
5 100.00 100.00 100.00 100.00
10 100.00 100.00 100.00 100.00
30 100.00 100.00 100.00 100.00
50 100.00 100.00 100.00 100.00
@@ A.1 IIA
10000 100.00 100.00 100.00 100.00
30000 100.00 100.00 100.00 100.00
50000 99.99 100.00 100.00 100.00
Table 10: Percentage IIA Violations Top Two

### v3

@@ 5.2 Pareto Unanimity (U)
A Pareto pair is defined to be a pair of alternatives where all voters prefer one alternative in the pair over the other.
Of the systems considered, only Group 1 systems violate U. Even for these systems, U is nearly always satisfied. In elections with at least 10 voters, all systems have less than a 1% probability of violating U. As the number of voters increases, the frequency of violations decreases because Pareto pairs become less likely. Table 4 shows the frequency with which systems violate U for elections with 3 voters.
Table 4: Percentage Unanimity Violations for 3 Voters (IAC)
Voting System Number of alternatives
@@
Plurality 10.92 22.35 40.91 62.69
Instant Runoff 10.92 22.35 40.91 62.69
Top Two 10.92 22.35 40.91 62.69
Anti-Plurality 10.92 22.27 40.76 61.96
Minimax 10.92 13.07 18.94 29.31
Nanson 10.92 7.58 20.27 20.64
Although percentage U violations are very low in most cases, they are common in elections with many alternatives and few voters. These trends are consistent with literature [29, 20].
The results for Instant Runoff, Plurality, and Top Two are identical. Data for all Group 1 systems is identical for three alternatives, but differ for four or more alternatives as noted in literature [20]. For larger number of alternatives, Nanson and Minimax perform the best, with less than 30% violations with six alternatives.
@@ 5.3 Transitivity (TR)
Pairwise Majority is the only system studied that violates transitivity. Table 5 shows the percentage TR violations for Pairwise Majority.
@@
    For a fixed number of voters, as the number of alternatives increases, transitivity violations increase.
@@
    For a fixed number of alternatives, as the number of voters increase, transitivity violations decrease and appear to converge. As the number of voters tends to infinity, the probability of violating transitivity approaches a fixed value [22]. For three alternative elections using IAC, that value is 6.25%, which is consistent with our findings.
@@
These results match the literature [20, 26, 22].  [^1] For elections with at least four alternatives, the limiting value for transitivity violations using IC in our study exceeds those in Niemi and Weisberg (1968) because Niemi and Weisberg examine failures to attain a Condorcet winner while we examine transitivity violations. Bottom cycles can account for the difference.
5.4 Joint Violation of Arrow’s Criteria
Simulations were run to find the percentage of times a voting system violates U, IIA, and TR. The other criterion (ND, UD) were not simulated. The joint-violation (JV) frequencies, defined as the percentage of times systems violated any one criteria, serve as an indication of fairness of a voting system.
Key Observations
  1. 1.
    Pairwise Majority is the least likely to violate all criteria jointly. This is due to the fact that transitivity violations are rarer than IIA violations.
  2. 2.
    Of the systems that satisfy transitivity, Ranked Pairs has the lowest JV frequencies followed by Baldwin, Borda Count, and Coombs.
  3. 3.
    For a fixed number of voters, as the number of alternatives increases, the JV frequencies increase for every system. For elections with 5 or more alternatives, JV frequencies are almost always 100%.
Table 6: Joint-Violation Frequencies for 100 voters (IAC)
Voting System Number of alternatives
3 4 5 6
Group 1 100.00 100.00 100.00 100.00
Copeland 85.50 100.00 100.00 100.00
Dowdall 96.84 100.00 100.00 100.00
Borda 83.85 99.94 100.00 100.00
Coombs 85.50 100.00 100.00 100.00
Baldwin 82.86 99.84 100.00 100.00
Ranked Pairs 75.66 98.67 100.00 100.00
Pairwise Majority 7.23 28.51 56.13 78.15
The conclusion for three-alternative elections is consistent with literature. Dougherty and Heckelman (2020) agrees that Pairwise Majority has the lowest JV frequencies. However, because differences in the IIA data, which is discussed in Appendix B, results for Copeland, Coombs, Dowdall, Borda, Ranked Pairs, and Baldwin do not match exactly. These systems have higher JV frequencies than in the 2020 study.
6 Conclusion
This study contributes to the enduring challenge in social choice theory of identifying voting systems that perform best with respect to Arrow’s axiomatic criteria. While Arrow’s impossibility theorem demonstrates that no voting system can simultaneously satisfy all fairness conditions under general preference domains, our results provide an empirical assessment of how often different systems fail in practice. Extending prior work by Dougherty and Heckelman (2020), our simulations incorporate larger electorates and a broader range of alternatives, offering a more comprehensive picture of violation frequencies across commonly used preference aggregation rules.
Our central finding - that Pairwise Majority exhibits the lowest joint-violation frequency among all tested systems -reinforces prior theoretical observations (Gehrlein, 1994; Niemi & Weisberg, 1968) and suggests that even within the constraints of Arrow’s theorem, some systems may better approximate normative ideals in practical settings. Importantly, we demonstrate that the likelihood of violating Arrow’s conditions increases as the number of alternatives increases, a trend consistent with both theoretical expectations and empirical findings (Nurmi, 1983; Börgers, 2010). These insights are especially relevant in contemporary democratic societies where elections increasingly involve crowded candidate fields and diverse voter preferences.
The implications of these results extend beyond theoretical interest. As public debate over electoral reform intensifies, a clearer understanding of the probabilistic fairness of voting systems is essential. While perfect fairness is unattainable, the empirical identification of systems with relatively lower joint violation frequencies, such as Pairwise Majority, can inform institutional design. Moreover, our enhanced methodology for detecting IIA violations offers more precise estimates, thereby addressing discrepancies in previous computational studies.
Our findings help show that while Arrow’s theorem remains an immutable constraint, its practical implications are context-dependent and, to some extent, quantifiable. Future work might extend this analysis to include strategic behavior, correlated preferences, or alternative fairness frameworks (e.g., monotonicity, Condorcet consistency), continuing the search for voting systems that, while not flawless, are less prone to failure in the eyes of theory and society alike.
7 Declarations
Data Availability Statement The raw data generated during this study is available upon request. All relevant processed data is included within this article and the appendix. The Python code is avaliable to run here: https://github.com/review123-coder/VotingSystemAnonymous.
Appendix A Data Tables
A.1 IIA
Table 7: Percentage IIA Violations Group 1
@@ A.1 IIA
Table 8: Percentage IIA Violations Borda
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
3 97.35 99.78 99.98 100.00
4 73.63 99.96 100.00 100.00
10 94.79 100.00 100.00 100.00
30 99.97 100.00 100.00 100.00
100 100.00 100.00 100.00 100.00
10000 100.00 100.00 100.00 100.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
3 89.08 99.08 99.96 100.00
4 56.66 99.81 100.00 100.00
10 74.29 99.90 100.00 100.00
30 84.52 99.87 100.00 100.00
100 84.14 99.93 100.00 100.00
10000 84.12 99.90 100.00 100.00
Table 9: Percentage IIA Violations Coombs
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
3 97.35 99.78 99.98 100.00
4 73.63 99.96 100.00 100.00
10 94.79 100.00 100.00 100.00
30 99.99 100.00 100.00 100.00
100 100.00 100.00 100.00 100.00
10000 100.00 100.00 100.00 100.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
3 89.08 99.08 99.96 100.00
4 56.66 99.81 100.00 100.00
10 74.29 99.91 100.00 100.00
30 87.40 100.00 100.00 100.00
100 85.70 100.00 100.00 100.00
10000 85.91 99.99 100.00 100.00
Table 10: Percentage IIA Violations Dowdall
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
3 97.35 99.78 99.98 100.00
4 99.52 99.96 100.00 100.00
10 99.79 100.00 100.00 100.00
@@
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
3 89.08 99.08 99.96 100.00
4 95.03 99.81 100.00 100.00
10 91.80 100.00 100.00 100.00
30 96.72 100.00 100.00 100.00
100 96.91 100.00 100.00 100.00
10000 96.93 100.00 100.00 100.00
Table 11: Percentage IIA Violations Copeland
@@
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
3 89.08 99.08 99.96 100.00
4 56.66 99.81 100.00 100.00
10 97.75 100.00 100.00 100.00
30 99.80 100.00 100.00 100.00
100 100.00 100.00 100.00 100.00
10000 99.98 100.00 100.00 100.00
Table 12: Percentage IIA Violations Baldwin
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
3 97.35 99.78 99.98 100.00
4 73.63 99.96 100.00 100.00
10 94.79 99.99 100.00 100.00
30 99.96 100.00 100.00 100.00
@@
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
3 89.08 99.08 99.96 100.00
4 56.66 99.81 100.00 100.00
10 74.29 99.76 100.00 100.00
30 83.23 99.76 100.00 100.00
100 83.35 99.75 100.00 100.00
10000 83.14 99.74 100.00 100.00
Table 13: Percentage IIA Violations Ranked Pairs
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
3 58.43 92.39 99.47 99.93
4 73.63 93.71 99.43 99.98
10 92.88 99.48 100.00 100.00
30 99.32 99.94 100.00 100.00
@@
10000 100.00 100.00 100.00 100.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
3 50.32 89.86 99.26 99.99
4 56.66 89.77 99.26 99.95
10 71.64 97.23 99.92 100.00
30 75.24 98.54 99.99 100.00
100 75.97 98.51 100.00 100.00
10000 75.87 98.66 100.00 100.00
A.2 Unanimity
Table 14: Percentage U Violations Plurality, Top Two, Instant Runoff
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
3 2.65 17.88 39.84 63.07
4 0.48 4.73 14.22 27.71
10 0.00 0.00 0.01 0.04
30 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
10000 0.00 0.00 0.00 0.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
3 10.92 22.35 40.91 62.69
4 4.97 7.87 15.37 28.00
10 0.22 0.06 0.01 0.08
30 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
Table 15: Percentage U Violations Minimax
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
3 2.65 9.25 18.01 29.83
4 0.48 1.71 3.73 6.68
10 0.00 0.00 0.00 0.00
30 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
10000 0.00 0.00 0.00 0.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
3 10.92 13.07 18.94 29.31
4 4.97 3.81 4.16 7.18
10 0.22 0.02 0.00 0.00
30 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
Table 16: Percentage U Violations Anti-Plurality
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
3 2.65 17.20 40.29 62.79
4 0.48 4.82 14.19 28.51
10 0.00 0.00 0.03 0.04
30 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
10000 0.00 0.00 0.00 0.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
3 10.92 22.27 40.76 61.96
4 4.97 7.98 15.12 28.12
10 0.22 0.03 0.02 0.04
30 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
Table 17: Percentage U Violations Nanson
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
3 2.65 5.09 19.27 21.60
4 0.48 2.70 5.63 9.50
10 0.00 0.00 0.00 0.01
30 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
10000 0.00 0.00 0.00 0.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
3 10.92 7.58 20.27 20.64
4 4.97 5.20 6.25 9.43
10 0.22 0.05 0.02 0.02
30 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
A.3 Joint Violation Frequencies
Table 18: JV Frequencies Group 1
(a) IC Data
Number of voters Number of alternatives
3 4 5 6